# Dealing with Uncertainty in Contextual Anomaly Detection

### v1

@@ Abstract
Contextual anomaly detection (CAD) aims to identify anomalies in a target (behavioral) variable conditioned on a set of contextual variables that influence the normalcy of the target variable but are not themselves indicators of anomaly. In this work, we propose a novel framework for CAD, normalcy score (NS), that explicitly models both the aleatoric and epistemic uncertainties. Built on heteroscedastic Gaussian process regression, our method regards the Z-score as a random variable, providing a contextual anomaly score but also a high-density interval that reflect the reliability of the contextual anomaly assessment. Through experiments on benchmark datasets and a real-world application in cardiology, we demonstrate that NS outperforms state-of-the-art CAD methods in both detection accuracy and interpretability. Moreover, high-density intervals enable an adaptive, uncertainty-driven decision-making process, which may be very important in domains such as healthcare.
@@ 1 Introduction
This particular problem setting is known as conditional (or contextual) anomaly detection (CAD) (Song et al. 2007; Valko et al. 2011; Tang et al. 2013; Liang and Parthasarathy 2016; Li and Van Leeuwen 2023). The name originates from the simple observation that rather than looking at the joint distribution of all variables, we split variables in two groups: contextual variables, $x$, and behavioral variables [^1] In (Song et al. 2007), contextual variables are called environmental attributes and behavioral variables are called indicator attributes., $y$. We then fit a conditional model and say that a given $x$ is anomalous if $p(y|x)$ is small according to our model. Note that modeling the conditional rather than the joint distribution is statistically more efficient, allowing us to work in application domains where data is not abundant.
@@
Related techniques started to be developed much earlier in medical statistics for constructing reference intervals (RIs) and centile curves (Goldstein 1972; Healy 1978; Royston 1991; Altman 1993; Zierk, Metzler, and Rauh 2021; Ammer et al. 2023). One example of application is the construction of child grow standards that provide percentiles of height and weight stratified by age and sex (WHO Multicentre Growth Reference Study Group 2006). In this body of literature (which, interestingly, does not intersect with the more recent machine learning CAD literature), contextual variables $x$ (such as age and sex) play the role of covariates while $y$ (e.g., height) plays the role of a response variable and is typically a continuous scalar. A classic approach is to assume that $y|x$ (perhaps after performing some variable transformations (Box and Cox 1964)) is normally distributed, fit a regression model to predict $y$ from $x$, and finally compute a Z-score (Roman et al. 1989; Lee et al. 2010; Colan 2013) as the difference between predictions and measurement, scaled by error standard deviation (see also Section 2). A Z-score above 2 is a typical threshold to flag an anomaly.
@@
A common trait of both CAD and methods for constructing RIs is that anomalous (out-of-reference) data points are identified by analyzing the conditional distribution $p(y|x)$. Effectively, these approaches are trying to model the aleatoric uncertainty (AU), which originates from the intrinsic variability of the behavior within the subpopulation that share the same context. The (implicit) underlying assumption is that anomalous data points are exactly those with large AU. From the point of view of the model, there is no way of reducing AU, except maybe by including more contextual variables (which may not even be possible). However, there is another source of uncertainty: we may lack samples in certain regions of the data space. This is known as epistemic uncertainty (EU) (Senge et al. 2014; Hüllermeier and Waegeman 2021), and in the present setting it occurs when a certain context is underrepresented in the data. On the one hand EU may arise because certain subpopulations are inherently (very) small and not included in the data. On the other hand, EU may arise due to various ethical, practical, and logistic challenges, which is the case in medical domains such as pediatrics (Ammer et al. 2023; Zierk, Metzler, and Rauh 2021; Ceriotti 2012), or due to various exclusion criteria, which is the case for imaging studies in cardiology (Ricci et al. 2021; Asch et al. 2019; Lancellotti et al. 2013). In our running example, overweight children may not appear in the data for these reasons. Accounting for the EU is crucial because normalcy assessment may be unreliable if the model’s EU is large.
@@
Despite the importance of modeling both AU and EU, to the best of our knowledge neither existing CAD nor normalcy approaches in medicine distinguish between them. This paper fills this gap. Specifically, we instantiate this idea as heteroscedastic Gaussian process regression (HGPR) (Goldberg, Williams, and Bishop 1997) that uses separate Gaussian processes to model the mean and standard deviation. This approaches confers several advantages and overcomes some well-known limitations of classic Z-scores (see, e.g., (Mawad et al. 2013; Colan 2013; Curtis et al. 2016)). First, it models the fact that the variablity of behavior depends on the context. Consequently, what is considered anomalous behaviour is coupled to the specific context. Second, using two Gaussian processes enables disentangling the AU in the data and EU in the model. This idea has been previously explored for active learning (Patel, Batra, and Murphy 2022) but not for anomaly detection. Third, being a Bayesian approach, it regards $Z$ as a random variable and thus naturally provides a method for computing high-density intervals on the anomaly score. This allows conveying a sense of the uncertainty of the model’s assigned anomaly score.
@@
  • We validate our method on standard publicly available benchmarks, consistently outperforming established CAD and AD methods in terms of ROC-AUC and PR-AUC.
@@ 2 Related Work
Figure 1: Simulated data from WHO growth curves for girls. Here $H\sim\mathcal{N}(\overline{h}(A),\overline{\sigma}(A))$, where the age-indexed mean function $\overline{h}(A)$ and standard deviation $\overline{\sigma}(A)$ are derived from published data (WHO Multicentre Growth Reference Study Group 2006). In the training data, $A\sim\mathrm{Exp(0.4)}$, so that young girls have a much higher probability of appearing in the training data. Non-contextual anomaly detection algorithms (such as isolation forests) trained to find the support of the joint $p(H,A)$ fail to spot children that are too tall or too short for their age, and additionally flag as outliers data points where the marginal $p(A)$ is small. ROCOD, a state-of-the-art algorithm for CAD, does not suffer from this problem but fails to correctly estimate conditional outliers in the presence of context-specific AU (heteroscedasticity). Our approach (NS) correctly identifies conditional outliers and, in addition, can abstain when EU (estimated as the width of the 95% HDI, $i(H,A)$, see Section 3.2) is high: In the plot, points with $i(H,A)>2$ are colored in blue with an intensity proportional to $i(H,A)$).
@@
The foundational work on CAD was introduced in (Song et al. 2007). In that paper, contextual and behavioral distributions were modeled independently using Gaussian mixture models and their dependencies were learned through mapping functions. Anomalies can be then identified as deviations from the learned dependencies. However, the approach may overstimate outlierness of objects with unusual contexts, as it would happen by modeling the joint $p(x,y)$ (see also Figure 1).
@@
ROCOD (Liang and Parthasarathy 2016) aims to mitigate the context sparsity problem by introducing a dual modeling approach that leverages both local and global perspectives to detect anomalies. The local model in ROCOD estimates the expected behavior of an object based on its contextual neighbors, which are identified as the objects most similar in terms of contextual attributes. In contrast, the global model adopts a holistic view of the dataset, learning the overall dependencies between contextual and behavioral variables through regression models. These two models are integrated via an adaptive weighting mechanism. ROCOD dynamically adjusts the relative importance of the local and global models based on the density of contextual neighbors for each object. When an object has a sufficient number of contextual neighbors, the local model is prioritized. Conversely, in sparse contexts, the global model dominates, ensuring robust anomaly detection across diverse scenarios.
@@
QCAD (Li and Van Leeuwen 2023) builds upon CAD and ROCOD by using quantile regression forests (Meinshausen and Ridgeway 2006) to model $p(y|x)$. Unlike ROCOD, which relies on the conditional mean to infer expected behavior, QCAD estimates conditional quantiles, providing a more detailed representation of the underlying data distribution. The anomaly score is computed by estimating the deviation of the observed $y$ from the predicted $p(y|x)$, looking at the width of the conditional percentile interval in which $y$ falls. Intuitively, if $y$ lies outside the estimated quantile range, the score is adjusted proportionally.
@@
Due to its simplicity and interpretability, the Z-score is widely used in normalcy studies (Roman et al. 1989; Colan 2013; Campens et al. 2014) and in the development of reference intervals (Lee et al. 2010; Zierk, Metzler, and Rauh 2021). In our context, the classic Z-score is defined as follows:
$Z(x,y)=\frac{y-f(x)}{S}$
@@
where $y$ and $x$ are the measured values of the behavioral variable and the contextual variables, respectively, $f$ is a model trained to predict behavior from context, and $S$ an estimation of the standard deviation (such as the square root of the square loss on the training set). Often, $f$ is a linear model, perhaps after the application of some ad-hoc Box-Cox transformations (Box and Cox 1964) (for example, one might regress the log of height on age in Figure 1 in order to model the saturation of growth).
@@
The above Z-score definition does not model heteroscedasticity in the data, i.e., situations when variability of the behavioral variable in the populations depends on the given context. Since the denominator averages out this variability, $Z(x,y)$ is overestimated in low-variance contexts and underestimated in high-variance contexts (Chubb and Simpson 2012; Mawad et al. 2013). Early approaches attempted to reduce heteroscedasticity by splitting data into separate groups (Goldstein 1972). Later, (Altman 1993) proposed a way to estimate variance by training a second linear regression model on the absolute residuals of the first model, a technique that may be seen as a simplified, linear, and non-Bayesian version of the technique proposed in this paper.
@@ 3 Methodology
This section describes our proposed normalcy score (NS) which tailors the classic Z-score to the contextual anomaly detection setting. Unlike the Z-score, NS is a random variable, even when the model and the test point $(x,y)$ are fixed. Like in (Goldberg, Williams, and Bishop 1997), we use two Gaussian process regression (GPR) models: one to model the mean and one to model the variance of the target variable as functions of the contextual features. We use GPR for three reasons. First, it provides a flexible non-parametric framework for modeling $p(y|x)$ as a Gaussian distribution. Intuitively, GPR defines a prior over functions $f(x)$ with variance $\sigma^{2}(x)$ using a mean function $m(x)$ and a covariance (kernel) function $k(x,x^{\prime})$, which encodes assumptions about smoothness and relationships between data points. Second, it allows disentangling the aleatoric and epistemic uncertainty. Finally, it permits providing both a point estimate and a high-density interval for an object’s anomaly score.
We begin by formalizing the definition of NS and its underlying Bayesian framework, then we describe how high-density intervals are derived for anomaly scores.
@@ 3.1 Normalcy Score
NS quantifies the deviation of a target variable $y$ from its expected behavior, relative to the contextual variance. To define NS, we first introduce two key components modeled using Gaussian process regression (GPR): $f_{1}$ and $f_{2}$.
@@
We model the log-standard deviation to ensure that the predicted standard deviation remains strictly positive following (Goldberg, Williams, and Bishop 1997).
@@
A key challenge with standard GPs is their running time, which scales as $O(N^{3})$, where $N$ is the number of points in the dataset. To address this, we use a sparse approximation based on inducing variables (Snelson and Ghahramani 2005; Titsias 2009). Using a set of $M$ inducing points to summarize the GP posterior and approximate the covariance matrix, the running time is reduced to $O(M^{2}N)$, where $M\ll N$.
We use NS to disentangles aleatoric uncertainty (AU) and epistemic uncertainty (EU) as follows:
@@ 3.2 Characterizing the NS distribution
Computing the high-density interval. One key aspect of the Bayesian approach is that it offers a principled method for quantifying the EU associated with the predicted mean and scale by looking at the concentration of the density $p(\mathrm{NS}(x,y))$. A simple way of expressing the concentration of $p(\mathrm{NS}(x,y))$ is by computing its highest density interval (HDI) (Kruschke 2015), defined as the smallest interval of values of NS such that $p(\mathrm{NS}(x,y))$ integrates to a given threshold (usually 95%). There is no general analytic form for the HDI. Hence, we first obtain samples of NS by sampling from $f_{1}$ and $f_{2}$ and then estimate $p$ by kernel density estimation. We finally perform numerical integration. We denote the length of the 95% interval as $i(x,y)$. As the dataset size increases, EU decreases and $p(\mathrm{NS}(x,y))$ becomes more concentrated, resulting in a smaller HDI.

### v2

@@ Abstract
Contextual anomaly detection (CAD) aims to identify anomalies in a target (behavioral) variable conditioned on a set of contextual variables that influence the normalcy of the target variable but are not themselves indicators of an anomaly. In this work, we propose a novel framework for CAD, normalcy score (NS), that explicitly models both the aleatoric and epistemic uncertainties. Built on heteroscedastic Gaussian process regression, our method regards the Z-score as a random variable, providing a contextual anomaly score but also a high-density interval that reflect the reliability of the contextual anomaly assessment. Through experiments on benchmark datasets and a real-world application in cardiology, we demonstrate that NS outperforms state-of-the-art CAD methods in both detection accuracy and interpretability. Moreover, high-density intervals enable an adaptive, uncertainty-driven decision-making process, which may be very important in domains such as healthcare.
@@ 1 Introduction
This particular problem setting is known as conditional (or contextual) anomaly detection (CAD) (Song et al., 2007; Valko et al., 2011; Tang et al., 2013; Liang and Parthasarathy, 2016; Li and Van Leeuwen, 2023). The name originates from the simple observation that rather than looking at the joint distribution of all variables, we split variables in two groups: contextual variables, $x$, and behavioral variables [^1] In Song et al. (2007), contextual variables are called environmental attributes and behavioral variables are called indicator attributes., $y$. We then fit a conditional model and say that a given $x$ is anomalous if $p(y|x)$ is small according to our model. Note that modeling the conditional rather than the joint distribution is statistically more efficient, allowing us to work in application domains where data is not abundant.
@@
Related techniques started to be developed much earlier in medical statistics for constructing reference intervals (RIs) and centile curves (Goldstein, 1972; Healy, 1978; Royston, 1991; Altman, 1993; Zierk et al., 2021; Ammer et al., 2023). One example application is the construction of child grow standards that provide percentiles of height and weight stratified by age and sex (WHO Multicentre Growth Reference Study Group, 2006). In this body of literature (which, interestingly, does not intersect with the more recent machine learning CAD literature), contextual variables $x$ (such as age and sex) play the role of covariates while $y$ (e.g., height) plays the role of a response variable and is typically a continuous scalar. A classic approach is to assume that $y|x$ (perhaps after performing some variable transformations (Box and Cox, 1964)) is normally distributed, fit a regression model to predict $y$ from $x$, and finally compute a Z-score (Roman et al., 1989; Lee et al., 2010; Colan, 2013) as the difference between predictions and measurement, scaled by error standard deviation (see also Section 2). A Z-score above 2 is a typical threshold to flag an anomaly.
@@
A common trait of both CAD and methods for constructing RIs is that anomalous (out-of-reference) data points are identified by analyzing the conditional distribution $p(y|x)$. Effectively, these approaches are trying to model the aleatoric uncertainty (AU), which originates from the intrinsic variability of the behavior within the subpopulation that share the same context. The (implicit) underlying assumption is that anomalous data points are exactly those with large AU. From the point of view of the model, there is no way of reducing AU, except maybe by including more contextual variables (which may not even be possible). However, there is another source of uncertainty: we may lack samples in certain regions of the data space. This is known as epistemic uncertainty (EU) (Senge et al., 2014; Depeweg et al., 2018; Hüllermeier and Waegeman, 2021), and in the present setting it occurs when a certain context is underrepresented in the data. On the one hand EU may arise because certain subpopulations are inherently (very) small and not included in the data. On the other hand, EU may arise due to various ethical, practical, and logistic challenges, which is the case in medical domains such as pediatrics (Ammer et al., 2023; Zierk et al., 2021; Ceriotti, 2012), or due to various exclusion criteria, which is the case for imaging studies in cardiology (Ricci et al., 2021; Asch et al., 2019; Lancellotti et al., 2013). In our running example, overweight children may not appear in the data for these reasons. Accounting for the EU is crucial because normalcy assessment may be unreliable if the model’s EU is large.
@@
Despite the importance of modeling both AU and EU, to the best of our knowledge neither existing CAD nor normalcy approaches in medicine distinguish between them. This paper fills this gap. Specifically, we instantiate this idea as heteroscedastic Gaussian process regression (HGPR) (Goldberg et al., 1997) that uses separate Gaussian processes to model the mean and standard deviation. This approach confers several advantages and overcomes some well-known limitations of classic Z-scores (see, e.g., (Mawad et al., 2013; Colan, 2013; Curtis et al., 2016)). First, it models the fact that the variability of behavior depends on the context. Consequently, what is considered anomalous behavior is coupled to the specific context. Second, using two Gaussian processes enables disentangling the AU in the data and EU in the model. This idea has been previously explored for active learning (Patel et al., 2022) but not for anomaly detection. Third, being a Bayesian approach, it regards $Z$ as a random variable and thus naturally provides a method for computing high-density intervals on the anomaly score. This allows conveying a sense of the uncertainty of the model’s assigned anomaly score.
@@
  • We validate our method on standard publicly available benchmarks, consistently outperforming established CAD and AD methods in terms of ROC AUC and PR AUC.
@@ 2 Related Work
Figure 1: Simulated data from WHO growth curves for girls. Here $H\sim\mathcal{N}(\overline{h}(A),\overline{\sigma}(A))$, where the age-indexed mean function $\overline{h}(A)$ and standard deviation $\overline{\sigma}(A)$ are derived from published data (WHO Multicentre Growth Reference Study Group, 2006). In the training data, $A\sim\mathrm{Exp(0.4)}$, so that young girls have a much higher probability of appearing in the training data. Non-contextual anomaly detection algorithms (such as isolation forests) trained to find the support of the joint $p(H,A)$ fail to spot children that are too tall or too short for their age, and additionally flag as outliers data points where the marginal $p(A)$ is small. ROCOD, a state-of-the-art algorithm for CAD, does not suffer from this problem but fails to correctly estimate conditional outliers in the presence of context-specific AU (heteroscedasticity). Our approach (NS) correctly identifies conditional outliers and, in addition, can abstain when EU (estimated as the width of the 95% HDI, $i(H,A)$, see Section 3.2) is high: In the plot, points with $i(H,A)>2$ are colored in blue with an intensity proportional to $i(H,A)$).
@@
The foundational work on CAD was introduced in Song et al. (2007). In that paper, contextual and behavioral distributions were modeled independently using Gaussian mixture models and their dependencies were learned through mapping functions. Anomalies can be then identified as deviations from the learned dependencies. However, the approach may overstimate outlierness of objects with unusual contexts, as it would happen by modeling the joint $p(x,y)$ (see also Figure 1).
@@
ROCOD (Liang and Parthasarathy, 2016) aims to mitigate the context sparsity problem by introducing a dual modeling approach that leverages both local and global perspectives to detect anomalies. The local model in ROCOD estimates the expected behavior of an object based on its contextual neighbors, which are identified as the objects most similar in terms of contextual attributes. In contrast, the global model adopts a holistic view of the dataset, learning the overall dependencies between contextual and behavioral variables through regression models. These two models are integrated via an adaptive weighting mechanism. ROCOD dynamically adjusts the relative importance of the local and global models based on the density of contextual neighbors for each object. When an object has a sufficient number of contextual neighbors, the local model is prioritized. Conversely, in sparse contexts, the global model dominates, ensuring robust anomaly detection across diverse scenarios.
@@
QCAD (Li and Van Leeuwen, 2023) builds upon CAD and ROCOD by using quantile regression forests (Meinshausen and Ridgeway, 2006) to model $p(y|x)$. Unlike ROCOD, which relies on the conditional mean to infer expected behavior, QCAD estimates conditional quantiles, providing a more detailed representation of the underlying data distribution. The anomaly score is computed by estimating the deviation of the observed $y$ from the predicted $p(y|x)$, looking at the width of the conditional percentile interval in which $y$ falls. Intuitively, if $y$ lies outside the estimated quantile range, the score is adjusted proportionally.
@@
Due to its simplicity and interpretability, the Z-score is widely used in normalcy studies (Roman et al., 1989; Colan, 2013; Campens et al., 2014) and in the development of reference intervals (Lee et al., 2010; Zierk et al., 2021). In our context, the classic Z-score is defined as follows:
$Z(x,y)=\frac{y-f(x)}{S}$ (1)
@@
where $y$ and $x$ are the measured values of the behavioral variable and the contextual variables, respectively, $f$ is a model trained to predict behavior from context, and $S$ an estimation of the standard deviation (such as the square root of the square loss on the training set). Often, $f$ is a linear model, perhaps after the application of some ad-hoc Box-Cox transformations (Box and Cox, 1964) (for example, one might regress the log of height on age in Figure 1 in order to model the saturation of growth).
@@
The above Z-score definition does not model heteroscedasticity in the data, i.e., situations when variability of the behavioral variable in the populations depends on the given context. Since the denominator averages out this variability, $Z(x,y)$ is overestimated in low-variance contexts and underestimated in high-variance contexts (Chubb and Simpson, 2012; Mawad et al., 2013). Early approaches attempted to reduce heteroscedasticity by splitting data into separate groups (Goldstein, 1972). Later, Altman (1993) proposed a way to estimate variance by training a second linear regression model on the absolute residuals of the first model, a technique that may be seen as a simplified, linear, and non-Bayesian version of the technique proposed in this paper.
@@ 3 Methodology
This section describes our proposed normalcy score (NS), which tailors the classic Z-score to the contextual anomaly detection setting. Unlike the Z-score, NS is a random variable, even when the model and the test point $(x,y)$ are fixed. As in Le et al. (2005), we jointly train two Gaussian process regression (GPR) models: one to model the mean and one to model the variance of the target variable as functions of the contextual features. We use GPR for three reasons. First, it provides a flexible non-parametric framework for modeling $p(y|x)$ as a Gaussian distribution. Intuitively, GPR defines a prior over functions $f(x)$ with variance $\sigma^{2}(x)$ using a mean function $m(x)$ and a covariance (kernel) function $k(x,x^{\prime})$, which encodes assumptions about smoothness and relationships between data points. Second, it allows disentangling the aleatoric and epistemic uncertainty. Finally, it permits providing both a point estimate and a high-density interval for an object’s anomaly score.
We begin by formalizing the definition of NS and its underlying Bayesian framework; then we describe how high-density intervals are derived for anomaly scores.
@@ 3.1 Normalcy Score
NS quantifies the deviation of a target variable $y$ from its expected behavior relative to the contextual variance. To define NS, we first introduce two key components modeled using Gaussian process regression (GPR): $f_{1}$ and $f_{2}$.
@@
We model the log-standard deviation to ensure that the predicted standard deviation remains strictly positive following Goldberg et al. (1997).
@@
A key challenge with standard GPs is their running time, which scales as $O(N^{3})$, where $N$ is the number of points in the dataset. To address this, we use a sparse approximation based on inducing variables (Snelson and Ghahramani, 2005; Titsias, 2009). Using a set of $M$ inducing points to summarize the GP posterior and approximate the covariance matrix, the running time is reduced to $O(M^{2}N)$, where $M\ll N$.
We use NS to disentangle aleatoric uncertainty (AU) and epistemic uncertainty (EU) as follows:
@@
To summarize:
  • We place independent GP priors on $f_{1}$ and $f_{2}$, modeling respectively the mean and the log standard deviation of $y\mid x$.
  • To make inference scalable, we adopt a sparse variational approximation with inducing points.
  • Training is performed using natural gradient descent for the variational parameters and the Adam optimizer for the remaining parameters.
@@ 3.2 Characterizing the NS distribution
Computing the high-density interval. One key aspect of the Bayesian approach is that it offers a principled method for quantifying the EU associated with the predicted mean and scale by looking at the concentration of the density $p(\mathrm{NS}(x,y))$. A simple way of expressing the concentration of $p(\mathrm{NS}(x,y))$ is by computing its highest density interval (HDI) (Kruschke, 2015), defined as the smallest interval of values of NS such that $p(\mathrm{NS}(x,y))$ integrates to a given threshold (usually 95%). There is no general analytic form for the HDI. Hence, we first obtain samples of NS by sampling from $f_{1}$ and $f_{2}$ and then estimate $p$ by kernel density estimation. We finally perform numerical integration. We denote the length of the 95% interval as $i(x,y)$. As the dataset size increases, EU decreases and $p(\mathrm{NS}(x,y))$ becomes more concentrated, resulting in a smaller HDI.
Aleatoric and epistemic uncertainty. As more data accumulate in the neighbourhood of $x$, the posterior means $m_{1}(x)$ and $m_{2}(x)$ converge to the underlying regression functions, while the posterior variances $\sigma_{1}^{2}(x)$ and $\sigma_{2}^{2}(x)$ shrink. This captures the reduction of epistemic uncertainty with increasing sample size. In contrast, the aleatoric component $e^{m_{2}(x)}$, representing the intrinsic variability of $y\mid x$, does not vanish even in the limit of infinite data. The expectation of $\mathrm{NS}(x,y)$ therefore contains a multiplicative correction factor $e^{\sigma_{2}^{2}(x)/2}$ arising from the log-normal distribution of $e^{-f_{2}(x)}$; this term converges to $1$ as $\sigma_{2}^{2}(x)$ decreases and ensures that the point estimate $s(x,y)$ is consistent with the exact posterior expectation. The length of the high-density interval $i(x,y)$, on the other hand, depends on both $\sigma_{1}^{2}(x)$ and $\sigma_{2}^{2}(x)$ and shrinks as these posterior variances decrease, making it a faithful indicator of epistemic uncertainty.